OpenStudy (anonymous):

Which is an example of the associative property? A. –(18 + 3) = –18 + (–3) B. –17 + 0 = –17 C. (–8 + 10) + 4. = –8 + (10 + 4) D. –5 + 7 = 7 + (–5)

OpenStudy (anonymous):

oh is it c?

Mehek (mehek14):

\(\tt{(a+b)+c=a+(b+c)}\) look for something following that rule

OpenStudy (anonymous):

yea... i think it is :/

OpenStudy (anonymous):

yea i put c :)

Mehek (mehek14):

yes C is correct ^_^

OpenStudy (anonymous):

thanks! :)

OpenStudy (anonymous):

could i ask you another? :)

Mehek (mehek14):

sure

OpenStudy (anonymous):

What is the value of the expression? 19.2 + (–7 + 3.8) A. 16 B. 24 C. 29 D. 30

Mehek (mehek14):

-7 + 3.8 = ?

OpenStudy (anonymous):

-3.2

OpenStudy (anonymous):

a?

Mehek (mehek14):

yes ^_^

OpenStudy (anonymous):

thanks! :)

Mehek (mehek14):

np :)

OpenStudy (anonymous):

What value of m makes the statement true? –(m + 15) = –18 + (–15) A. –33 B. –18 C. 3 D. 18

Mehek (mehek14):

\(\bf{-m-15=-18-15\\-m=-18\\m=18}\)

OpenStudy (anonymous):

thank you again! :p

OpenStudy (anonymous):

What value of x makes the equation true? –(9 + (–17)) = –9 + x

OpenStudy (anonymous):

how r u so good at this stuff? XD

OpenStudy (anonymous):

@pooja195